Neptune Company produces toys and other items for use in beach and resort areas. A small, inflatable toy has come onto the market that the company is anxious to produce and sell. The new toy will sell for $3.40 per unit. Enough capacity exists in the company’s plant to produce 30,700 units of the toy each month. Variable expenses to manufacture and sell one unit would be $2.14, and fixed expenses associated with the toy would total $57,523 per month.

The company's Marketing Department predicts that demand for the new toy will exceed the 30,700 units that the company is able to produce. Additional manufacturing space can be rented from another company at a fixed expense of $2,876 per month. Variable expenses in the rented facility would total $2.38 per unit, due to somewhat less efficient operations than in the main plant.

Required:

1. What is the monthly break-even point for the new toy in unit sales and dollar sales.

2. How many units must be sold each month to attain a target profit of $12,342 per month?

3. If the sales manager receives a bonus of 15 cents for each unit sold in excess of the break-even point, how many units must be sold each month to attain a target profit that equals a 22% return on the monthly investment in fixed expenses?

(For all requirements, Round "per unit" to 2 decimal places, intermediate and final answers to the nearest whole number.)
Answer #1

1.

breakeven units sales=total fixed cost/contribution per unit

total fixed cost of month= fixed expenses associated with the toy + fixed expense of rent per month.

total fixed cost=$57523+$2876=$60399

contribution per unit=selling price per unit - variable manufacturing cost per unit

selling price per unit=$3.40

variable manufacturing cost per unit=$2.14

contribution per unit=$3.40-$2.14=$1.26

breakeven units sales=$60399/$1.26=47935.71 rounding to the nearest whole number 47936 units

breakeven units sales in dollar sales= breakeven units sales*selling price

breakeven units sales in dollar sales= 47936*3.40=$162982.4

2)

to attain a target profit of $12,342 per month

targeted sales in dollar=(breakeven sales in dollar sales + targeted profit)/contribution per unit

targeted sales in dollar=($162982.4+$12324)/$1.26=$139132.06

targetes asales in units=targeted sales in dollar/selling price per unit

targetes asales in units=$139132.06/$3.40=40921.19 rounding to the nearest whole number 40921 units

3)

for every sales after breakeven poin manager recieves a bonus of 15 cents for each unit sold and to attain 22% return on the monthly investment in fixed expenses

total fixed expense in month= $57523+$2876=$60399

22% return on fixed expense=$60399*22%=$13287.78

number of units to be sold after breakeven point to attain the targeted profit of 22% return on fixed expense =22% return on fixed expense/15 cents

15 cents=$0.15

units must be sold each month to attain a target profit that equals a 22% return on the monthly investment in fixed expenses =$13287.78/0.15=88585.2 rounding to the nearest whole number 88585 units

units must be sold each month to attain a target profit that equals a 22% return on the monthly investment in fixed expenses = breakeven unit sale + units must be sold each month to attain a target profit that equals a 22% return on the monthly investment in fixed expenses

units must be sold each month to attain a target profit that equals a 22% return on the monthly investment in fixed expenses = 47936+88585=136521 units